when salsa cost $3 there are 400 bags of chips purchased when salsa cost $4 there are 150 bags of chips purchased what is the CPED of chips-salsa at those salsa price levels
Answer #1

Answer


the cross-price elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=150-400=-250
average quantity=(150+400)/2=275
change in price=4-3=1
average price=(4+3)/2=3.5

cross-price elasticity of demand=(-250/275)/(1/3.5)
=-3.18181818
=-3.19
The elasticity is -3.19 so the two goods are complements as the price and quantity are inversely related.